# Can the 3 neutrino masses really be found using SN 1987A data?

### v2

@@ IV.7 Suppose neutrinos all have $m<<1eV/c^{2}$?
It is natural to wonder what Fig. 1 would look like if the three neutrino masses were very small and none were tachyonic, as in the standard neutrino description. Clearly, they still would lie on three lines, since our derivation leading to the conclusion of three straight lines on a $1/E^{2},t$ plot just depended on relativistic kinematics. However, we almost certainly would not be able to find these lines because for 3 small masses we would find three nearby steeply sloped lines, as in Fig.5. showing how the 3 lines might appear for neutrinos emitted at $t=0$ and $t=1$ sec Of course, with a horizontal error bar of 1 or 2 sec, identifying which of 30 points went with which nearly vertical line would be impossible because there are in fact many more line-triplets between $t=0$ and $t=1sec.$
If the neutrino mass were in fact under the KATRIN current upper limit $(0.45eV/c^{)}$ the data points would all lie on lines that were virtually vertical and the data points would all lie in the shaded central area in Fig.1 having a 2 second time window. This is obviously far from reality, which suggests that their mass is not below the KATRIN limit. In the “double bang” scenario for SN 1987A which could conceivably explain the LSD burst, the preceding remains true, so once again we would not be able to identify the three lines about which the 30 points are clustering, unlike the actual data in Fig. 1. This means that the double bang scenario can be ruled out, given Fig.1.
Figure 5: The case of neutrinos having very small masses (which yield triplets of steeply sloped straight lines nearly overlapping) with a data point having a 2 sec wide horizontal error bar. Neutrinos are assumed emitted at $t=0$ and $t=1$ seconds. Given the size of the error bar, it would be impossible to find which of the six near vertical lines the data point is associated with (likewise for all the other line triplets for $0>t>1$ seconds – completely unlike what is actually the case in Fig.1.
IV.8 Consistency with direct mass experiments?
@@
The three masses deduced from the three straight line fit in Fig. 1 seem preposterously large, and they would appear to be inconsistent with data, whether it be cosmological data or data from the KATRIN experiment implying $m_{\nu}<0.45eV/c^{2}$. As previously discussed, KATRIN measures an effective mass, and if one of the three mass states contributing to the electron neutrino effective mass has $m_{k}^{2}<0$ the effective mass, as normally defined, $m_{e}^{2}=\sum|U_{e,k}|^{2}m_{k}^{2}$ can be arbitrarily close to zero or even negative, even when the $m_{k}^{2}$ are all very large, as in the SN 1987 fit. KATRIN has further crucial relevance to that fit quite apart from the limit it imposes by its measured $\nu_{e}$ effective mass.
IV.9 A missing $m_{4}=2.70eV/c^{2}$ sterile neutrino?
@@
IV.10 Looking for the wrong kind of neutrinos?
A likely possibility is that KATRIN failed to see any $m>1eV/c^{2}$ neutrinos because they were looking for the wrong kind of neutrinos in their sterile neutrino search. They should have been looking for active $m>1eV/c^{2}$ neutrinos not sterile ones – based on what was found in our SN 1987A data analysis. Any $m>1eV/c^{2}$ active neutrinos unfortunately would not show up given the way the KATRIN sterile search was conducted. Thus, obviously, KATRIN should not have included in the fitted combined differential spectrum any zero mass contribution. Rather they needed to fit their data using a combined differential spectrum $R_{C}$ for.the three masses $m_{1}=2.70eV/c^{2},m_{2}=22.4eV/c^{2}$ and $m_{3}^{2}=-200,000eV^{2}/c^{4}$ with three weights, $\alpha_{j},$ that is:
@@ IV.10 Looking for the wrong kind of neutrinos?
To test the validity of the main claim of this paper (that the three mass fit of SN 1987A data yields the three true masses), KATRIN could see where on the $\alpha_{1}$ and $\alpha_{2}$ plane good fits to the spectrum are found – doing this both for the three stipulated masses and away from them. Finally, for the functional form of the spectrum for the tachyonic third mass, it may be possible to use the usual beta decay formula, and simply substitute a negative $m^{2}$ in it. However, Ciborowski and Rembielinski have shown that for their tachyonic neutrino model a slightly more complicated formula is needed whose specific form depends on whether helicity or chirality coupling is used – see Eqs. 20-21 in ref. (Ciborowski1999, )
@@
If KATRIN were to redo its analysis with these suggested changes it would be a decisive test of whether the SN 1989A data really can provide the three neutrino mass values, including the tachyonic one $m_{3}$. A positive KATRIN result which confirms all three masses from the SN 1987A data fit would be of some considerable interest, while a negative result would doom the idea of obtaining neutrino masses from the SN 1987A data. The missing $2.70eV/c^{2}$ sterile neutrinos in KATRIN was the key clue here, in the proverbial “dog that didn’t bark.” category.
IV.11 Consistency with cosmological data?
In the recent literature the neutrino mass sum is considered to be the sum over mass state masses. However, going back to its original usage by Gershtein and Zeldovitch in 1966, the sum of the neutrino masses was intended to be a constraint on the sum of the flavor states – then only $\nu_{e}$ and $\nu_{\mu}.$ (Gershtein1966, ) However, at that time no distinction was made between mass and flavor states. Apart from historical usage there is a reason to view the sum as being over flavor state masses. Cosmological neutrinos were probably born in the universe’s infancy, decoupling from matter and radiation when the universe was only about 1 second old. At that time the two dominant processes capable of producing neutrinos was $Z^{0}$ decay or $e^{+}e^{-}$ collisions into neutrino antineutrino pairs, and would both produce equal numbers for each flavor. For example, $Z^{0}$ decays would result in equal $6.7\%$ branching ratios for each flavor, thus creating an equal number density N of each. Likewise for $e^{+}e^{-}$ collisions. In this case, the cosmological neutrino mass/energy density would be:
$R=Nm_{\nu,e}+Nm_{\nu,\mu}+Nm_{\nu,\tau}=N\Sigma m_{\nu}$ (7)
so, the appropriate constraint from Eq. 7 should be on the flavor state sum, and not the mass state sum, as usually assumed. Thus, even though the three mass state $m^{2}$ values from the SN 1987A fit are all very large, as long as one of them has $m^{2}<0,$ the three flavor state effective masses can all be very tiny, which appears to be the case.
@@ V Summary and the ultimate test
This paper has made six claims that some readers may consider very difficult to believe – see Table 1. Some of the claims will likely not be accepted as true without the strongest corroborating evidence. After KATRIN does the analysis suggested in section IV.9, and it finds evidence for the three active neutrino masses deduced from the fit to SN 1987 data with little uncertainty in their values, that might be enough.
@@ 4. One of the 3 $m^{2}_{j}$ from SN 1987A data is a $m^{2}<0$ tachyon
6. KATRIN likely can prove or reject all the above
@@
However, the most convincing test would be a new galactic supernova. Given the much larger neutrino detectors today compared to those in 1987, a new galactic supernova might yield thousands of neutrinos reaching the detectors. This 88 year-old does not have a great chance of seeing it, but let’s be an optimist. When supernova SN 2026(!) occurs at a distance that is a percentage p of the distance of SN 1987A, and we were to find an LSD-like near-monochromatic burst of a thousand neutrinos occurring earlier than the main neutrino burst by a time interval of $484\times p$ minutes, there would be little doubt that tachyons exist, as well as the other five claims. Even if the next galactic supernova does not occur in 2026, and I don’t get to see it, it is comforting to realize the evidence, in the form of thousands of neutrinos are now on their way to Earthly detectors. These neutrinos will definitely either confirm or reject the seven claims made. Moreover, given the known frequency of supernovae, we know that those neutrinos are now more than $99.99\%$ of their way here from some star in our galaxy. No one knows which star it will be, but Betelgeuse, a massive red supergiant in Orion, is expected to end its life in a spectacular supernova explosion soon. Soon in astronomical terms means potentially within the next 100,000 years (some say under 300 years (Saio2023, )). Were Betelgeuse to become a supernova, it would become incredibly bright, and easily visible in daylight. Go Betelgeuse!
@@ Acknowledgments
The author is grateful to Alan Chodos and Donald Gelman for helpful comments. If KATRIN performs the decisive test suggested in section IV.9 while I am still on this planet, I thank Guido Drexlin and Diana Parno of the KATRIN Group.

### v3

@@ IV.7 Suppose neutrinos have $m<<1eV/c^{2}$?
It is natural to wonder what Fig. 1 would look like if the three neutrino masses were small and none were tachyonic, as in the standard neutrino description. Clearly, they still would lie on three lines if emitted simultaneously, since our derivation leading to the conclusion of three straight lines on a $1/E^{2},t$ plot just depended on relativistic kinematics. However, we would not be able to find these lines because for 3 small masses they would lie on very many almost vertical lines filling the time window – see the 2 second error bar in Fig, 5. In effect each point lies on its own vertical line filling the time window. This is obviously far from what is the case for the actual data, which suggests that the neutrino mass does not satusfy $m<<1eV/c^{2}.$
Figure 5: How 50 neutrinos might appear on a $1/E^{2}$ versus t plot if their mass $m<<1eV/c^{2}$ and they were emitted during a time window of 2 seconds. No congregation of the points about specific lines would be observed, completely unlike the real SN 1987A data which clearly clusters along three straight lines. Essentially, each of the 50 points lies on its own different vertical line somewhere within the 2 second time window and no two points lie on the same line.
IV.8 Consistency with cosmological data?
In the recent literature the issue of whether the sum is over mass or flavor states is not widely discussed. However, going back to its original usage by Gershtein and Zeldovitch in 1966, the sum of the neutrino masses was intended to be a constraint on the sum of the flavor states – then only $\nu_{e}$ and $\nu_{\mu}.$ (Gershtein1966, ) However, at that time no distinction was made between mass and flavor states. Apart from historical usage there is a reason to view the sum as being over flavor state masses. Cosmological neutrinos were probably born in the universe’s infancy, decoupling from matter and radiation when the universe was only about 1 second old. At that time the two dominant processes capable of producing neutrinos was $Z^{0}$ decay or $e^{+}e^{-}$ collisions into neutrino antineutrino pairs, and would both produce equal numbers for each flavor. For example, $Z^{0}$ decays would result in equal $6.7\%$ branching ratios for each flavor, thus creating an equal number density N of each. Likewise for $e^{+}e^{-}$ collisions. Thus, the cosmological neutrino mass/energy density produced in that manner would be:
$R=Nm_{\nu,e}+Nm_{\nu,\mu}+Nm_{\nu,\tau}=N\Sigma m_{\nu}$ (6)
so, the appropriate constraint on R from fitting cosmological data is from Eq. 6 a constraint on the flavor state sum, and not the mass state sum, as sometimes assumed. Thus, even though the three mass state $m^{2}$ values from the SN 1987A fit are all very large, as long as one of them has $m^{2}<0,$ the three flavor state effective masses can all be very tiny, which appears to be the case.
IV.9 Consistency with direct mass experiments?
@@
The three masses deduced from the three straight line fit in Fig. 1 seem preposterously large, and they would appear to be inconsistent with data from the KATRIN experiment implying $m_{\nu}<0.45eV/c^{2}$. As previously discussed, KATRIN measures an effective mass, and if one of the three mass states contributing to the electron neutrino effective mass has $m_{k}^{2}<0$ the effective mass, as normally defined, $m_{e}^{2}=\sum|U_{e,k}|^{2}m_{k}^{2}$ can be arbitrarily close to zero or even negative, even when the $m_{k}^{2}$ are all very large, as in the SN 1987 fit. KATRIN has further crucial relevance quite apart from the limit it imposes by its measured $\nu_{e}$ effective mass.
IV.10 A missing $m_{4}=2.70eV/c^{2}$ sterile neutrino?
@@
IV.11 “The dog that didn’t bark” (the key section!)
It is possible that KATRIN failed to see any sterile neutrinos because they simply do not exist, despite the above high confidence kevel, but here we consider a much more interesting possibility. Neutrino - 4 and the other three experiments that saw the $2.70eV/c^{2}$ neutrino were doing a short baseline disappearance oscillation experiment, which observed the disappearance probability $P(\bar{\nu_{e}}\rightarrow\bar{\nu_{e}}).$ This type of experiment would work just as well if the disappearing neutrinos were disappearing into sterile or active neutrinos. However, KATRIN’s search was specifically looking for sterile neutrinos and it would not work for active ones as conducted. (KATRIN was not being careless here, they just assumed like everyone else that neutrinos with large mass would have to be sterile) Of course, if the main claim of this paper is correct, and if the $2.70eV/c^{2}$ neutrino is an active neutrino (as well as a sterile one - see Fig. 4), then KATRIN’s failure to see it is understandable because their search method just would not work for active neutrinos. Specifically, KATRIN should not have included in the fitted combined differential spectrum any zero mass contribution. (Why not you ask? Because the zero mass effective mass is made up of the three specific masses we found including one with $m^{2}<0$) Rather, they needed to fit their data using a combined differential spectrum $R_{C}$ for.the three masses $m_{1}=2.70eV/c^{2},m_{2}=22.4eV/c^{2}$ and $m_{3}^{2}=-200,000eV^{2}/c^{4}$ with three unknown adjustible weights, $\alpha_{j},$ that is:
@@
To test the validity of the main claim of this paper (that the three mass fit of SN 1987A data yields the three true masses), KATRIN could see where on the $\alpha_{1}$ and $\alpha_{2}$ plane good fits to the observed spectrum are found. With the three masses set to their specified SN 1987A values, KATRIN should observe a small concentrated area where good fits are found surrounded by a large area of poor fits. They could also do the same using other masses, where no good fits should be found. The best fit on the $\alpha_{1}$ and $\alpha_{2}$ plane should yield the effective mass from $m^{2}=\sum\alpha_{j}m^{2}_{j},$ and it should be a very small negative value.
Finally, for the functional form of the spectrum for the tachyonic third mass, $m_{3}^{2},$ they could just use the usual beta decay formula, and simply substitute a negative $m^{2}$ in it. While Rembielinski, Caban and Ciborowski have shown that their causal tachyonic neutrino model results in a more complicated formula – see Eqs. 93-94 in ref. (Rembielinski2021, ), however, as shown in their Fig. 1 the simple approach gives a result that is virtually indistinguishable from their full description.
@@
In summary, if KATRIN were to redo its analysis with the preceding suggested changes it could be a decisive test of whether the SN 1989A data really can provide the three neutrino mass values, including the tachyonic one $m_{3}$. A positive KATRIN result which confirms all three masses from the SN 1987A data fit would be of some considerable interest. The missing $2.70eV/c^{2}$ sterile neutrinos in the KATRIN search was the key clue here. It was effectively “the dog that didn’t bark,” a crucial example of a “negative fact.”
IV.12 Finding more evidence in oscillation data?
Neutrino oscillation experiments of the dissapearing type can also be used to find evidence of the three neutrino masses from the SN 1987A data using a similar approach. The key quantity fitted to the data in such an experiment is the survival probability $P_{j}$ of a neutrino of a given flavor. One usually looks to see if there is any significant amplitude $\alpha_{j}$ for the probability anywhere on a grid of values over sterile neutrino mass and amplitude of the oscillation for a single sterile neutrino.
To find how much amplitude oscillation there is in the data for neutrinos of mass $m_{1}=2.70eV/c^{2}$ and simultaneously for neutrinos of mass $m_{2}=22.4eV/c^{2}$ one can create a grid of $\alpha_{1}$ and $\alpha_{2}$ values with the summed probability $P_{1}+P_{2}$ calculated at each grid point, and see if there is a small region of the grid where the summed probabilities is very high. If so, that would constitute strong support for the idea that the $m^{2}$ masses of electron neutrinos consist of the values deduced from the SN 1987A data. The most promising experiments to look at might be the four identified in ref. (Serebrov2024, ) which already have shown evidence of the $2.70eV/c^{2}$ mass – do they now show evidence of a $22.4eV/c^{2}$ neutrino as well? Of course, only some experiments having a very short baseline will be sensitive to a mass this high. It would also be interesting to look at some formerly negative results such as one from the NOvA Collaboration to see if they see anything. (Acero2025, )
@@ V Summary and the ultimate test
This paper has made six claims that many. readers may understandably consider very difficult to believe – see Table 1. Some of the claims will likely not be accepted as true without the strongest corroborating evidence. After KATRIN does the analysis suggested in section IV.10, and it were to find evidence for the three active neutrino masses deduced from the fit to SN 1987 data with little uncertainty in their values, that might be enough. This might also be the case if dissapearance experiments were to find such evidence doing the analysis mentioned in the previous section.
@@ 4. One of the 3 $m^{2}_{j}$ from SN 1987A data is a $m^{2}<0$ tachyon
6. Existing data likely can prove or reject all the above
@@
However, the most convincing test would be a new galactic supernova. Given the much larger neutrino detectors today compared to those in 1987, a new galactic supernova might yield thousands of neutrinos reaching the detectors. This 88 year-old does not have a great chance of seeing it, but let’s be an optimist. When supernova SN 2026(!) occurs at a distance that is a percentage p of the distance of SN 1987A, and we were to find an LSD-like near-monochromatic burst of a thousand neutrinos occurring earlier than the main neutrino burst by a time interval of $484\times p$ minutes, there would be little doubt that tachyons exist, as well as the other five claims. Even if the next galactic supernova does not occur in 2026, and I don’t get to see it, it is comforting to realize the evidence, in the form of thousands of neutrinos are now on their way to Earthly detectors. These neutrinos will definitely either confirm or reject the six claims made. Moreover, given the known frequency of supernovae, we know that those neutrinos are now more than $99.99\%$ of their way here from some star in our galaxy. No one knows which star it will be, but Betelgeuse, a massive red supergiant in Orion, is expected to end its life in a spectacular supernova explosion soon. Soon in astronomical terms means potentially within the next 100,000 years (some say under 300 years (Saio2023, )). Were Betelgeuse to become a supernova, it would become incredibly bright, potentially as bright as the full moon. Go Betelgeuse!
@@ Acknowledgments
The author is grateful to Alan Chodos and Donald Gelman for helpful comments. I also thank Guido Drexlin and Diana Parno of the KATRIN Collaboration for their kindness.